# Towards Next-Generation Recommender Systems: A Benchmark for Personalized Recommendation Assistant with LLMs

### v1

@@ Abstract.
Recommender systems (RecSys) are widely used across various modern digital platforms and have garnered significant attention. Traditional recommender systems usually focus only on fixed and simple recommendation scenarios, making it difficult to generalize to new and unseen recommendation tasks in an interactive paradigm. Recently, the advancement of large language models (LLMs) has revolutionized the foundational architecture of RecSys, driving their evolution into more intelligent and interactive personalized recommendation assistants. However, most existing studies rely on fixed task-specific prompt templates to generate recommendations and evaluate the performance of personalized assistants, which limits the comprehensive assessments of their capabilities. This is because commonly used datasets lack high-quality textual user queries that reflect real-world recommendation scenarios, making them unsuitable for evaluating LLM-based personalized recommendation assistants. To address this gap, we introduce RecBench+, a new dataset benchmark designed to access LLMs’ ability to handle intricate user recommendation needs in the era of LLMs. RecBench+ encompasses a diverse set of queries that span both hard conditions and soft preferences, with varying difficulty levels. We evaluated commonly used LLMs on RecBench+ and uncovered below findings: 1) LLMs demonstrate preliminary abilities to act as recommendation assistants, 2) LLMs are better at handling queries with explicitly stated conditions, while facing challenges with queries that require reasoning or contain misleading information. Our dataset has been released at https://github.com/jiani-huang/RecBench.git.
@@
copyright: acmlicensedjournalyear: 2018isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Figure 1. Illustration of Different Recommendation Paradigms. (a) Traidtional-based recommendations generate suggestions in simple and fixed scenarios by learning user/item representations, like Top-K or similar item recommendations. (b) LLM-based recommendation relies on fixed and simple query templates, restricting users to predefined formats. (c) In contrast, our benchmark evaluates the LLM’s ability to handle complex, flexible user queries as a recommendation assistant.
@@ 1. Introduction
Despite the remarkable success, current studies lack a comprehensive evaluation of the recommendation capability of these LLM-based personalized recommendation assistants. Specifically, in the phases of both training and testing, most existing methods leverage fixed and simple task-specific prompt templates to generate recommendation responses, and assess the recommendation performance of these assistants, such as “Will the user like {movie_i}. Please answer Yes or No.” to determine whether to recommend a specific item, or “Please choose a movie from the candidate list {movie_1, movie_2, … }” to recommend a user-liked item. However, as shown in Figure 1 (c), this testing paradigm for evaluation fails to align with real-world practical scenarios, where the user queries can be highly complex when they interact with the LLM-based recommendation assistant, such as “Recommend me a funny romantic movie like Alien, starring Leonardo Wilhelm DiCaprio, and suitable for female audiences.” In this case, the recommendation assistant must not only understand the user’s desire to find movies starring Leonardo Wilhelm DiCaprio, but also filter out funny romantic movies that are appropriate for female viewers. Moreover, the recommendation assistant should be robust to perturbations in user input, i.e., it should recognize that the user is looking for a funny animated movie rather than a horror film like Alien. This requires the assistant to deeply comprehend multiple constraints within the user’s input and provide suitable and robust recommendations through remarkable reasoning capability, which is quite challenging. However, promising results on the existing benchmark datasets may not be extended to the new recommendation paradigm in the era of LLMs.
@@
Existing frequently used datasets, such as LastFM (Schedl, 2016), Amazon Beauty (Ni et al., 2019), and Movielens-1M (Harper and Konstan, 2015), typically only include basic user information, item metadata (e.g., title, categories, and descriptions), and user-item interactions. These datasets are more suitable for assessing the accuracy performance of traditional RecSys but are not appropriate for effectively evaluating recommendation assistants in the era of LLMs, due to the lack of high-quality textual user queries which can meet practical recommendation scenarios.
@@
To bridge this gap, we introduce a new benchmark, RecBench+, tailored for evaluating the potential of LLMs to act as personalized recommendation assistants. Specifically, we propose a new dataset comprising approximately 30,000 high-quality and complex user queries, constructed based on multiple frequently-used recommendation datasets (e.g., Movielens-1M (Harper and Konstan, 2015) and Amazon book [^1] https://archive.ics.uci.edu/ml/datasets/Amazon+book+reviews). These queries vary in difficulty, number of conditions that user specified, and user profiles, effectively reflecting the diverse user needs in real-world scenarios. We comprehensively evaluate the performance of seven widely used LLMs on RecBench+, including GPT-4o (Hurst et al., 2024), Gemini-1.5-Pro (Team et al., 2024), DeepSeek-R1 (Guo et al., 2025), etc. We further analyze various factors that impact model performance, such as the number of conditions and the integration of user interaction history. Through extensive experiments with state-of-the-art LLMs in our benchmark, we obtain several insightful findings: (1) LLMs demonstrate preliminary abilities to act as recommendation assistant, but their strengths vary. GPT-4o and DeepSeek-R1 excel at user queries with clear conditions, while Gemini-1.5-Pro and DeepSeek-R1 perform better in queries without clear conditions but require a clear understanding of user profiles. (2) For queries with user-specified conditions, LLMs are more adept at handling those with explicitly stated conditions. In contrast, queries that require reasoning or contain perturbations pose greater challenges. However, models equipped with advanced reasoning capabilities tend to perform better on such tasks. (3) For queries requiring user profile understanding, LLMs show varying performance across different user interests and demographics. For example, LLMs may be better at providing recommendations for popular interests and tend to perform more effectively for female users than for male users. Our contributions are summarized as follows:
@@
  • Novel Recommendation Paradigm: We introduce a new paradigm for next-generation recommender systems, exploring the effectiveness of LLMs acting as personalized recommendation assistants in an interactive and intelligent way. As the first benchmark to effectively integrate LLMs in this interactive and intelligent manner, we transform the traditional recommendations paradigm into a more personalized and context-aware experience, setting a new standard for the field.
@@
  • Dataset Construction: We present a comprehensive dataset called RecBench+ to simulate practical recommendation scenarios in the era of LLMs. Our dataset encompasses approximately 30,000 high-quality and complex user queries that vary in difficulty, the number of conditions specified by users, and user profiles, reflecting the diverse user needs in real-world recommendation scenarios. To the best of our knowledge, this is the first public dataset that can be used to effectively evaluate the performance of personalized recommendation assistants in the era of LLMs.
@@
  • Comprehensive Evaluation: We conduct extensive experiments with state-of-the-art LLMs, analyzing their strengths and limitations in this task and providing actionable insights.
@@ 2.1. Overview
In practical applications, user demands can generally be categorized into two types, each requiring specialized capabilities from LLM-based recommendation assistants. The first type is driven by specific conditions or constraints, such as asking for movies by a particular director or within a certain genre. This scenario demands strong reasoning capabilities from the recommendation assistant, as it requires the assistant to accurately interpret and apply the given constraints to generate appropriate recommendations. The second type does not involve specified conditions but requires recommendation assistants to generate recommendations based on an understanding of the user’s profile (e.g., user interest, occupation, gender), which emphasizes the assistant’s ability to interpret the profile and provide personalized suggestions. Given that these two scenarios demand different capabilities from the recommendation assistant, we construct two distinct query types to evaluate their performance: Condition-based and User Profile-based. Below, we introduce the two query categories in detail.
@@ 2.1.1. Condition-based Query
Condition-based Query simulates scenarios where users have a clear understanding of what they want and set concrete conditions for the expected items. For example, in a movie recommendation assistant, a user might seek films featuring specific directors, actors, genres, or other attributes. The core challenge here lies in correctly understanding and satisfying all constraints in the query, particularly when queries include multiple conditions. To comprehensively evaluate the recommendation assistant’s reasoning ability to fully interpret and fulfill constraints, we categorize Condition-based Query into three types according to how users express their needs when seeking recommendations. Typically, users explicitly state their conditions for the desired item, which we define as Explicit Condition Query. In addition to clearly stating the conditions, users may only provide vague or incomplete information about the items they need, requiring the recommendation assistant to infer the underlying conditions from the context, which we regard as Implicit Condition Query. Additionally, users might have mistaken information or make typos for the item, leading to Misinformed Condition Query. The following are detailed explanations:
  • Explicit Condition Query:
    These queries explicitly specify the conditions that the recommended items should satisfy in the natural language format. For example, the query, “Can you recommend movies featuring Gwyneth Paltrow?”, directly states the condition (“featured by Gwyneth Paltrow”) the recommendations must satisfy. Explicit condition queries are relatively straightforward because the constraints are clearly articulated.
  • Implicit Condition Query: Unlike Explicit Condition Query, Implicit Condition Query require the recommendation assistant to infer the conditions of the target items, because the user may not provide complete or specific details about the items. For example, “Can you suggest movies with the same cinematographer as Stay Hungry (1976) and Beethoven (1992)?” does not mention the cinematographer’s name (i.e., David Worth), forcing the recommendation assistant to conduct in-depth reasoning about the name of the cinematographer first. Implicit condition queries aim to assess assistants’ reasoning and knowledge retrieval capabilities, making them more challenging than explicit condition queries.
  • Misinformed Condition Query: Users may sometimes provide incorrect or misleading information about items in their queries. For example, the query, “I like ‘Avatar’ directed by Spielberg. Please recommend similar movies directed by him.”, incorrectly attributes the movie Avatar to Steven Spielberg. It is expected that recommendation assistants can first identify and correct this misinformation–recognizing that Avatar was not directed by Spielberg—before generating relevant recommendations. This subcategory presents a significant challenge to the robustness of the recommendation assistant, requiring it to accurately identify and filter out inaccurate information in users’ queries.
@@ 2.1.2. User Profile-based Query
  • Demographics-based Query: These queries focus on how users’ demographic attributes influence recommendations. Different users, based on their age, gender, or occupation, may have distinct preferences and needs. For example, philosophy professors might frequently be interested in films that explore moral dilemmas. By leveraging these patterns, we can craft queries such as: “I’m a philosophy professor looking for movies with complex themes and moral dilemmas. Can you recommend any?” By aligning with real-world demographic behaviors, these queries can mirror the personalized recommendations that a personalized recommendation assistant should provide, ensuring that the suggestions are relevant and tailored to the user’s background and identity.
2.1.3. Statistics of the Constructed Benchmark Dataset RecBench+
The benchmark dataset’s statistical breakdown is shown in Table 1, illustrating the distribution of query types. The movie dataset is sourced from Movielens-1M [^2] MovieLens-1M provides detailed user demographic information that can be used for demographics-based queries, while other versions, such as MovieLens-10M and MovieLens-20M, lack such information. Therefore, we only can adopt the MovieLens-1M dataset to construct our dataset., while the book dataset is from Amazon-Book. We then describe the construction details of Condition-based Query and User Profile-based Query.
@@ 2.1.3. Statistics of the Constructed Benchmark Dataset RecBench+
Number of Users 6,036 4,421
Number of Items 3,247 9,016
Number of User-Item Interactions 33,291 29,285
@@ 2.2. Condition-based Query Construction
Figure 2. An overview of the pipeline of constructing Condition-based Query. The process starts by constructing an Item Knowledge Graph (KG) that links items to their attributes. The second step is to identify item sets from the user’s history that share the same relations (e.g., movies directed by the same person) using the KG. These shared relations serve as the basis for constructing different types of conditions for the query. Finally, we use LLMs to simulate users and generate queries based on conditions, and the aforementioned item set serves as the ground truth.
@@ 2.2.3. Query Generation
    where the set $\{i_{1},i_{2},\ldots,i_{m}\}$ contains items from the KG that do not actually share the relationship $r_{i}$ with $t_{i}$. For example, given a condition (director, Cameron), a misinformed condition might be (director, Cameron, error info: Cameron is the director of ‘Startreck’).
@@
    where $\mathcal{P}$ is the prompt, see appendix H. For each generated query $q$, we construct a sample consisting of:
@@ 2.3. User Profile-based Query Construction
Figure 3. An overview of the pipeline of constructing User Profile-based Query. (a) For Interest-based Query (top), we identify common interests by extracting collaborative item sets and preceding item sequences from user interactions, and queries are generated based on inferred interests. (b) For Demographics-based Query (bottom), users are grouped by demographics, and queries are generated using demographics and popular items within each group.
In addition to providing specific constraints, users may provide only basic personal information such as watching history or demographic attributes (e.g., age, occupation). To evaluate the LLM recommendation assistant’s ability to handle such scenarios, we introduce User Profile-based Query, which aims to generate personalized recommendations based on user profile information without requiring specific filtering conditions. As detailed in Section 2.1.2, User Profile-based Query contains two subcategories: Interest-based Query and Demographics-based Query. The construction process is illustrated in Figure 3 and we describe the construction process below.
@@ 2.3.1. Interest-based Query
These queries are designed to capture user interests that are not explicitly defined by specific conditions but are inferred from user interactions. To effectively simulate real-world scenarios, we focus on capturing the collective interests that emerge from shared behaviors from multiple users, which are more representative of general preferences and trends. For example, if many users who watched “The Dark Knight” also watched “Inception”, this suggests a clear interest in action and suspense. In this way, we can craft queries that reflect collective user interests derived from shared behaviors. Specifically, we first identify Common Interest Set, which are sets of items that frequently interact consecutively by multiple users. Let $\mathcal{H}_{u}$ denote the interaction history of user $u$, and =$\mathcal{H}=\{\mathcal{H}_{u}\mid u\in\mathcal{U}\}$ denote the set of all users’ interaction histories. The common interest set can be defined as:
@@
Based on the identified common interest set and preceding item sequences, we employ an LLM to analyze the patterns and infer the reasoning behind multiple users’ common interests. The generated query aims to reflect both the contextual features of the items and the common interests of users. The prompt used to generate these queries is provided in Appendix H.
2.3.2. Demographics-based Query
Individuals with similar demographic backgrounds, like age, gender, or occupation, often exhibit similar preferences and interaction patterns. For example, kids typically enjoy animated movies like Disney films, while teenagers may prefer action and superhero movies. Therefore, by grouping users who share the same demographic attributes (e.g., students aged 18–24), we can effectively construct recommendation queries influenced by demographic attributes. Specifically, we categorize users into different user groups based on permutations and combinations of demographic attributes. For each user group, we first identify the set of items most frequently consumed by users within that group. Subsequently, we provide both the user group demographics and the list of most popular items as input to an LLM. The LLM then analyzes the underlying patterns of each user group. Subsequently, LLM generates a query that encapsulates both the distinct characteristics of the user group demographics and the preferences reflected in the most popular items. This query is designed to test the ability of recommendation assistants to provide personalized suggestions based on demographic attributes. The list of most popular items will serve as the ground truth for evaluating the relevance of recommendations generated in response to the query.
3. Experiments
3.1. Experiments Settings
3.1.1. Baselines
We evaluate several widely used models, including GPT-4o (2024-08-06) (Hurst et al., 2024), GPT-4o-mini (2024-07-18) (Hurst et al., 2024), Llama (Llama-3.1-70B-Instruct) (Dubey et al., 2024), Gemini (gemini-1.5-pro-002) (Team et al., 2023), Claude (claude-3-5-sonnet-20241022), DeepSeek-V3 (Liu et al., 2024a), and DeepSeek-R1 (Guo et al., 2025). However, we were unable to test the GPT-o1 (Jaech et al., 2024) model from OpenAI because it flagged our prompts as violating its usage policy.
3.1.2. Evaluation Metrics
We utilize the below metrics to evaluate LLM performance. To simulate real-world scenarios where users typically do not predefine the number of recommendations they need, we do not specify a fixed number of recommendations K in the testing prompts; however, we also include several experiments with a fixed K in the Appendix F for reference.
Table 2. Performance on Condition-based Query. The  /  means the best/the second-best result. In Misinformed Condition Query, a higher FTR rate is better as it shows the model can reject recommendations when queries have factual errors.

### v2

@@ Abstract.
Recommender systems (RecSys) are widely used across various modern digital platforms and have garnered significant attention. Traditional recommender systems usually focus only on fixed and simple recommendation scenarios, making it difficult to generalize to new and unseen recommendation tasks in an interactive paradigm. Recently, the advancement of large language models (LLMs) has revolutionized the foundational architecture of RecSys, driving their evolution into more intelligent and interactive personalized recommendation assistants. However, most existing studies rely on fixed task-specific prompt templates to generate recommendations and evaluate the performance of personalized assistants, which limits the comprehensive assessments of their capabilities. This is because commonly used datasets lack high-quality textual user queries that reflect real-world recommendation scenarios, making them unsuitable for evaluating LLM-based personalized recommendation assistants. To address this gap, we introduce RecBench+, a new dataset benchmark designed to assess LLMs’ ability to handle intricate user recommendation needs in the era of LLMs. RecBench+ encompasses a diverse set of queries that span both hard conditions and soft preferences, with varying difficulty levels. We evaluated commonly used LLMs on RecBench+ and uncovered below findings: 1) LLMs demonstrate preliminary abilities to act as recommendation assistants, 2) LLMs are better at handling queries with explicitly stated conditions, while facing challenges with queries that require reasoning or contain misleading information. Our dataset has been released at https://github.com/jiani-huang/RecBenchPlus.
@@
journalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USA.booktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAisbn: 979-8-4007-2292-9/2026/02doi: 10.1145/3773966.3777954ccs: Information systems Personalizationcopyright: none
Figure 1. Illustration of different recommendation paradigms. (a) Traditional-based: Generating recommendations in fixed scenarios by learning user/item representations. (b) LLM-based recommendation Employing fixed and simple query templates, lacking evaluation in more complex scenarios. (c) Ours: Evaluating the LLM’s ability to handle complex, flexible user queries as a recommendation assistant.
@@
Despite the remarkable success, current studies lack a comprehensive evaluation of the recommendation capability of these LLM-based personalized recommendation assistants. Specifically, in the phases of both training and testing, most existing methods leverage fixed and simple task-specific prompt templates to generate recommendation responses, and assess the recommendation performance of these assistants, such as “Will the user like {movie_i}. Please answer Yes or No.” to determine whether to recommend a specific item, or “Please choose a movie from the candidate list {movie_1, movie_2, … }” to recommend a user-liked item. However, this evaluation paradigm fails to align with real-world practical scenarios, where the user queries can be highly complex when they interact with the LLM-based recommendation assistant. Consider the complex query illustrated in Figure 1 (c), the LLM-based recommendation assistant must first interpret the user’s description to identify the correct movie, then recognize the lead actor, and finally suggest similar films featuring that actor. This process requires the assistant to understand the query, perform multi-step reasoning, and consider multiple conditions—capabilities that go beyond basic attribute matching or shallow semantic understanding. Therefore, the results of existing LLM-based RecSys on traditional benchmark datasets may not apply to the new recommendation paradigm in the LLM era.
@@
Existing frequently used datasets, such as LastFM (Schedl, 2016), Amazon Beauty (Ni et al., 2019), and Movielens-1M (Harper and Konstan, 2015), typically only include basic user information, item metadata (e.g., title), and user-item interactions. These datasets are more suitable for assessing the accuracy performance of traditional RecSys but are not appropriate for effectively evaluating recommendation assistants in the era of LLMs, due to the lack of diverse complex user queries that can meet practical recommendation scenarios. To bridge this gap, we introduce a new benchmark, RecBench+, tailored for evaluating the potential of LLMs to act as personalized recommendation assistants. Specifically, we propose a new dataset of approximately 30,000 complex user queries, covering a range of recommendation scenarios that require complex reasoning, including multiple conditions, multi-hop reasoning, misleading information, and diverse user profiles. This dataset is built from several commonly used recommendation datasets (e.g., Movielens-1M (Harper and Konstan, 2015) and Amazon book [^1] https://archive.ics.uci.edu/ml/datasets/Amazon+book+reviews).
@@
We comprehensively evaluate the performance of seven widely used LLMs on RecBench+, including GPT-4o (Hurst et al., 2024), Gemini-1.5-Pro (Team et al., 2024), DeepSeek-R1 (Guo et al., 2025), etc. We further analyze various factors that impact model performance, such as the number of conditions, the integration of user interaction history and different fine-tuning paradigms. Through extensive experiments with state-of-the-art LLMs in our benchmark, we obtain several insightful findings: (1) LLMs demonstrate preliminary abilities to act as recommendation assistant, but their strengths vary. GPT-4o and DeepSeek-R1 excel at user queries with clear conditions, while Gemini-1.5-Pro and DeepSeek-R1 perform better in queries without clear conditions but require a clear understanding of user profiles. (2) Fine-tuning notably boosts LLM recommendation performance, with a two-stage approach combining supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT) performing best overall. This suggests that SFT effectively warms up the model for the task, enabling RFT to further refine performance. (3) For queries with user-specified conditions, LLMs are more adept at handling those with explicitly stated conditions. In contrast, queries that require reasoning or contain perturbations pose greater challenges. However, models equipped with advanced reasoning capabilities tend to perform better on such tasks. (4) For queries requiring user profile understanding, LLMs show varying performance across different user interests and demographics. For example, LLMs may be better at providing recommendations for popular interests and tend to perform more effectively for female users than for male users. Our contributions are summarized as follows:
@@
  • Novel Recommendation Paradigm: We introduce a new paradigm for next-generation recommender systems, exploring the effectiveness of LLMs acting as personalized recommendation assistants in an interactive and intelligent way. As the first benchmark to integrate LLMs in this way, we transform the traditional recommendation paradigm into a more personalized and context-aware experience, setting a new standard for this field.
@@
  • Dataset Construction: We present a comprehensive dataset called RecBench+ to simulate practical recommendation scenarios in the era of LLMs. Our dataset encompasses approximately 30,000 complex user queries that require complex reasoning, including multiple conditions, multi-hop reasoning, misleading information, and diverse user profiles. To the best of our knowledge, this is the first public dataset that can be used to evaluate the performance of LLM-based recommendation assistants in complex and reasoning-intensive recommendation settings.
@@
  • Comprehensive Evaluation: We conduct extensive experiments with state-of-the-art LLMs and fine-tune open-source models for this task, analyzing their strengths and limitations in this task and providing actionable insights.
@@ 1.1. Overview
In practical applications, user demands can generally be categorized into two types, each requiring specialized capabilities from LLM-based recommendation assistants. The first type is driven by specific conditions or constraints, such as asking for movies by a particular director or within a certain genre. This scenario demands strong reasoning capabilities from the recommendation assistant, as it requires the assistant to accurately interpret and apply the given constraints to generate appropriate recommendations. The second type does not involve specified conditions but requires recommendation assistants to generate recommendations based on an understanding of the user’s profile (e.g., user interest, occupation, gender), which emphasizes the assistant’s ability to interpret the profile and provide personalized suggestions. Given that these two scenarios demand different capabilities from the recommendation assistant, we construct two distinct query types to evaluate their performance: Condition-based and User Profile-based. As shown in Table 1, our benchmark utilizes Amazon-Book and Movielens-1M [^2] Unlike the 10M/20M versions, MovieLens-1M includes essential user demographic data.. Below, we detail the construction of Condition-based and User Profile-based queries based on these sources.
@@ 1.1.1. Condition-based Query
The Condition-based Query simulates scenarios where users impose concrete constraints (e.g., specific directors or genres). The core challenge lies in accurately satisfying these conditions. To evaluate reasoning capabilities, we categorize this query into three types: Explicit Condition Query (directly stated constraints), Implicit Condition Query (vague inputs requiring multi-hop reasoning), and Misinformed Condition Query (containing errors or typos).
  • Explicit Condition Query: These queries explicitly specify the conditions that the recommended items should satisfy in the natural language format. For example, the query, “Can you recommend movies featuring Gwyneth Paltrow?”, directly states the condition (“featured by Gwyneth Paltrow”) the recommendations must satisfy. Explicit condition queries are relatively straightforward because the constraints are clearly articulated.
  • Implicit Condition Query: Unlike Explicit Condition Query, Implicit Condition Query involves incomplete specifications, demanding multi-hop reasoning to deduce target conditions. For example, requesting “movies with the same cinematographer as Stay Hungry and Beethoven” requires inferring the unstated name (David Worth) first. This category serves as a more challenging benchmark for reasoning capabilities.
  • Misinformed Condition Query: Users may sometimes provide incorrect or misleading information about items in their queries. For example, the query, “I like ‘Avatar’ directed by Spielberg. Please recommend similar movies directed by him.”, incorrectly attributes the movie Avatar to Steven Spielberg. It is expected that recommendation assistants can first identify and correct this misinformation–recognizing that Avatar was not directed by Spielberg—before generating relevant recommendations. This subcategory presents a significant challenge to the robustness of the recommendation assistant, requiring it to accurately identify inaccurate information in users’ queries.
@@ 1.1.2. User Profile-based Query
Number of Users 6,036 48,535
Number of Items 3,247 50,088
Number of User-Item Interactions 29,459 55,424
Figure 2. Overview of the pipeline of constructing Condition-based Query.
  • Demographics-based Query: These queries focus on how users’ demographic attributes influence recommendations. Different users, based on their age, gender, or occupation, may have distinct preferences and needs. For example, philosophy professors might be interested in films that explore moral dilemmas. By leveraging these patterns, we can craft queries such as: “I’m a philosophy professor looking for movies with complex themes and moral dilemmas. Can you recommend any?” By aligning with real-world demographic behaviors, these queries can mirror the personalized recommendations that a personalized recommendation assistant should provide, ensuring that the suggestions are relevant and tailored to the user’s background and identity.
@@ 1.2.2. Shared Relation Extraction
Figure 3. Overview of the pipeline of constructing User Profile-based Query.
@@ 1.2.3. Query Generation
    where the set $\{i_{1},i_{2},\ldots,i_{m}\}$ contains items from the KG that do not actually share the relationship $r_{i}$ with $t_{i}$. For example, a misinformed condition can be (director, Cameron, error info: Cameron is the director of ‘Startreck’).
@@
    where $\mathcal{P}$ is the prompt. For each generated query $q$, we construct a sample consisting of:
@@ 1.3. User Profile-based Query Construction
In addition to providing specific constraints, users may provide only basic personal information such as watching history or demographic attributes (e.g., age, occupation). To evaluate the LLM recommendation assistant’s ability to handle such scenarios, we introduce User Profile-based Query, which aims to generate personalized recommendations based on user profile information without requiring specific filtering conditions. As detailed in Section 1.1.2, User Profile-based Query contains two subcategories: Interest-based Query and Demographics-based Query. The construction process is illustrated in Figure 3 .
@@ 1.3.1. Interest-based Query
These queries are designed to capture user interests that are not explicitly defined by specific conditions but are inferred from user interactions. To effectively simulate real-world scenarios, we focus on capturing the collective interests that emerge from shared behaviors from multiple users, which are more representative of general preferences and trends. For example, if many users who watched “The Dark Knight” also watched “Inception”, this suggests a clear interest in action and suspense. In this way, we can craft queries that reflect collective user interests derived from shared behaviors. Specifically, we first identify Common Interest Set, which are sets of items that frequently interact consecutively by multiple users. Let $\mathcal{H}_{u}$ denote the interaction history of user $u$, and $\mathcal{H}=\{\mathcal{H}_{u}\mid u\in\mathcal{U}\}$ denote the set of all users’ interaction histories. The common interest set can be defined as:
@@
Table 2. Performance on Condition-based Query. The  /  means the best/the second-best result.
@@
Based on the identified common interest set and preceding item sequences, we employ an LLM to analyze the patterns and infer the reasoning behind multiple users’ common interests. The generated query aims to reflect both the contextual features of the items and the common interests of users.
1.3.2. Demographics-based Query
Individuals with similar demographics often share preferences; for instance, children typically favor animated films (e.g., Disney), whereas teenagers may prefer action or superhero movies. Leveraging this, we group users by attribute permutations (e.g., students aged 18–24) and identify their top-consumed items. We then input these demographics and item lists into an LLM to synthesize queries that capture group-specific patterns. The identified popular items serve as the ground truth for evaluation.
1.4. Data Filtering and Selection
For Condition-based Query, we use string matching to check if the provided information (e.g., actor, director) appears in the query and filter out queries that omit this information. For User Profile-based Query, we manually verify the queries and filter out those that are too broad or unreasonable.
2. Experiments
2.1. Experiments Settings
2.1.1. Baselines
We evaluate several widely used models, including GPT-4o (2024-08-06) (Hurst et al., 2024), GPT-4o-mini (2024-07-18) (Hurst et al., 2024), Llama (Llama-3.1-70B-Instruct) (Dubey et al., 2024), Gemini (gemini-1.5-pro-002) (Team et al., 2023), Claude (claude-3-5-sonnet-20241022), DeepSeek-V3 (Liu et al., 2024a), and DeepSeek-R1 (Guo et al., 2025) on our proposed dataset. However, we were unable to test the GPT-o1 (Jaech et al., 2024) model because it flagged our prompts as violating its usage policy. In addition, we further perform full-parameter supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT, based on the GRPO (Shao et al., 2024) algorithm) on the Qwen-2.5-3B-Instruct and LLaMA-3.2-3B-Instruct models to explore their potential performance gains under task-specific optimization.
2.1.2. Evaluation Metrics
We utilize the below metrics to evaluate LLM performance. To simulate real-world scenarios where users typically do not predefine the number of recommendations they need, we do not specify a fixed number of recommendations K in the testing prompts; however, we also include several experiments with a fixed K in Section 2.2 for reference.